OpenStudy (anonymous):

What is the surface area of a cone with a 6 in. diameter and a length (slant height) of 8 in.?

geerky42 (geerky42):

SA = πr² + πrs where r represents radius and s represent slant height. = π(3)² + π(3)(8) = 9π + 24π = (9+24)π = 33π

OpenStudy (anonymous):

which one? A) 11.6 in.2 B) 75.4 in.2 C) 104 in.2 D) 179 in.2

geerky42 (geerky42):

You can just make an approximate guess, we know that π is close to 3 so 33·3 = 99, which is close to 104 so C.

OpenStudy (anonymous):

thanks a lot :3